### v2

@@ 1.1 Relevance
A well-designed intervention often consists of more than one phase. This is because, to prevent fadeouts, the benefit of a treatment in an earlier phase may need to be reinforced through follow-up treatments in the later phases (Cunha & Heckman, 2007). This rationale underpins many sustained treatment strategies in fields such as education, job training, counseling, and medicine for promoting long-term human development and well-being. In evaluating such programs, a research quantity of key scientific interest is the cumulative average treatment effect (ATE) of an intended experimental treatment sequence compared to a control sequence for a target population.
Authorized by the Tennessee state legislature in the 1980s, the Student Teacher Achievement Ratio study (Project STAR) was a state-wide multisite randomized trial designed to evaluate the multi-year cumulative impact of class size reduction on student achievement. Upon kindergarten entry, students within each participating school were randomly assigned to either a small class of 13-17 students (i.e., the experimental group) or a regular class of 22-25 students (i.e., the control group). Students were expected to remain in the class size type initially assigned throughout the study years. At the beginning of each year, teachers within each school were also randomly assigned by the project staff to teach a small class or a regular class (Finn and Achilles, 1990; Mosteller et al., 1995). In this two-year study (kindergarten followed by Grade 1), each year represents a separate phase as students are assessed at the end of a year and are expected to possibly join a different class and meet a different teacher in the coming year. Our application study examines the cumulative ATE of receiving two years of instruction in a small class compared to a regular class in both kindergarten and Grade 1 in student achievement at the end of Grade 1.
@@ 1.2 Methodological Challenges
Noncompliance following the Phase I treatment. Noncompliance often becomes inevitable when an intervention program involves successive phases. In Project STAR, the annual progression in schooling naturally provides an occasion, at the end of each school year, for the school to move some students from one treatment condition to another, even though such decisions might be contrary to the intended research design. Indeed, some students changed their class size types after the kindergarten year (Hanushek, 1999). Identifying the cumulative ATE is challenging in the presence of noncompliance due to preexisting measured and unmeasured differences between those who complied with the initial treatment assignment and those who did not.
Posttreatment confounding. If an individual does not respond well to a Phase I treatment, it may seem sensible to adjust their treatment in the subsequent phase when possible. In the context of class size interventions, for example, reasons for changing a student’s class size type in Grade 1 are likely tied to the student’s performance in kindergarten. However, kindergarten performance is a post-Phase 1 treatment (but pre-Phase II treatment) covariate that may confound the causal relationship between the Grade 1 treatment receipt and the outcome. Importantly, regression-based adjustments for such posttreatment confounding can introduce bias when assessing the effect of the kindergarten (Phase I) treatment on the outcome (Rosenbaum, 1984).
Non-additive treatment effects. In some cases, continuing the same treatment into a later phase may not only sustain but also amplify the impact of the Phase I treatment. As a result, the benefit of receiving the treatment in two consecutive phases may exceed the sum of the benefits of receiving the treatment in Phase I alone and Phase II alone. In some other cases, a second phase of the same treatment may offer little additional value, such that the combined benefit of receiving two phases of the treatment is less than the sum of the single-phase treatment effects. For these reasons, past studies that examined the impact of class size reduction in either kindergarten or Grade 1 in isolation have not addressed the key question about the cumulative ATE of receiving the treatment across both years.
@@ 1.3 Constraints of Existing Causal Inference Strategies
Inverse-probability-of-treatment weighting (IPTW) is commonly used to address posttreatment confounding when evaluating the cumulative effects of time-varying treatments (Hong and Raudenbush, 2008; Robins et al., 2000) or per-protocol effects in the presence of nonadherence (Hernán and Robins, 2017; Smith et al., 2021). However, identification using IPTW relies on the sequential ignorability assumption that rarely holds in practice, as noncompliant behavior is often related to unobserved confounders, even after adjusting for observed ones. In prior analyses of the Project STAR data, researchers have employed statistical adjustment for observed pretreatment covariates but failed to account for both posttreatment confounding and unobserved pretreatment confounding (Nye et al., 2000).
@@
When applying a standard instrumental variable (IV) strategy to address noncompliance, the analyst may use the initial random treatment assignment as the instrument for the non-random treatment received in a later phase. The IV estimand is defined as the ratio of the ITT effect of the instrument on the outcome to the ITT effect on the treatment received (Bloom, 1984). This strategy requires the exclusion restriction, which is unlikely to hold in multi-phase interventions. This is because the instrument is expected to have a direct effect on the outcome through actual exposure to the Phase I treatment, regardless of compliance in Phase II. Therefore, the standard IV method is not appropriate for identifying the cumulative ATE in such settings (Swanson et al., 2018). Previous studies using Project STAR data have applied the IV method to reduce selection bias associated with noncompliance in year-by-year analyses. However, these approaches do not address the cumulative ATE, as the yearly effects may not be additive (Krueger, 1999; Shin and Raudenbush, 2011).
@@
Researchers have sought to estimate the average effect of the Phase II treatment on the outcome in the target population by extrapolating from the complier population. This is done by first estimating compliance scores as functions of observed pretreatment covariates that predict both compliant behavior and heterogeneity in treatment effects, and then estimating the average effect of the Phase II treatment within levels of compliance scores (Angrist and Fernandez-Val, 2010; Aronow and Carnegie, 2013; Esterling et al., 2011; Wang et al., 2017). These strategies nonetheless rely on a form of the strong ignorability assumption that is often unrealistic. Specifically, unobserved pretreatment confounders as well as observed and unobserved time-varying confounders may not be captured by the estimated compliance scores, which undermines the validity of the identification.
@@
The multiple-site multiple-mediator IV (MSMM-IV) strategy (Bloom et al., 2020; Duncan et al., 2011; Kling et al., 2007; Nomi et al., 2021; Raudenbush et al., 2012; Reardon and Raudenbush, 2013), developed to identify the causal effects of multiple parallel mediators in a multisite trial, similarly requires the exclusion restriction and overlooks non-additive cumulative treatment effects. This strategy generally does not allow for time-varying confounders of the mediator-outcome relationships that may affect the outcome both directly and indirectly (see Reardon and Raudenbush (2013) for a discussion of a possible relaxation of this constraint).
@@ 1.4 Proposed Multisite Two-Phase Treatment IV Strategy
When using the initial random treatment assignment as an instrument to address complications arising from noncompliance, the well-known principal stratification framework (Frangakis and Rubin, 2002) assumes deterministic noncompliance. That is, each individual is classified as either a complier or a noncomplier, based on the notion that compliance is pre-determined and fixed. Under this framework, the IV estimand identifies the Complier Average Treatment Effect (CATE), also known as the Local Average Treatment Effect (LATE) (Angrist et al., 1996).
However, as Small et al. (2017) argue, this deterministic view may not align with reality, especially when compliant behavior is influenced by stochastic events. In such cases, the IV estimand instead identifies a Weighted Average Treatment Effect (WATE), a quantity that often does not directly correspond to the target population of the study. In our application, we similarly conceptualize noncompliance as stochastic. For the primary target population in Project STAR, we argue that the concepts of CATE or LATE may not be applicable while the concept of cumulative ATE of attending a small versus regular class over consecutive years is relevant.
@@ 2.1 Notation and the Causal Estimand
Let $Z_{ik}$ be a binary indicator for the Phase I treatment that takes value $z=1$ if student $i$ in school $k$ is assigned to a small class and 0 if assigned to a regular class upon kindergarten entry. Due to full compliance in kindergarten, the treatment assigned is equivalent to the treatment received for each student in Phase I. Let $D_{ik}$ be a binary indicator for the Phase II treatment receipt that takes value $d=1$ if the student attends a small class and 0 if attending a regular class in Grade 1. The student’s potential achievement by the end of Grade 1 can be represented as a function of the initial randomized treatment assignment $z$ or a function of the two-year treatment sequence $z$ and $d$. It is typically denoted as $Y_{ik}(z)$ in the former case and $Y_{ik}(z,d)$ in the latter case for $z=0,1$ and $d=0,1$ under the stable unit treatment value assumption (SUTVA). In other words, we assume that there is a single version of each treatment and that there is no interference among students between or within schools. Similarly, we may use $V_{ik}(z)$ to denote the student’s potential achievement by the end of kindergarten and use $D_{ik}(z)$ to denote the student’s Grade 1 treatment receipt each as a function of the student’s kindergarten treatment assignment. We use $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$ to denote the student’s observed and unobserved vectors of baseline covariates, respectively. In this multisite randomized trial, every participating school is a site. We use $S_{i}=k$ to indicate that student $i$ is a member of school $k$.
Corresponding to a research question that initially motivated Project STAR, the causal effect of attending a small class versus a regular class in both kindergarten and Grade 1 for student $i$ in school $k$ is $Y_{ik}(1,1)-Y_{ik}(0,0)$. The cumulative ATE, as defined below, represents the average of school-specific treatment effects for a population of schools where the inner expectation is taken over the population of students within school $k$ and the outer expectation is taken over all the schools:
@@ 2.2 Theoretical Models
Figure 1 illustrates the conceptual relationships between $X,U,Z,V,D$, and $Y$. We have omitted the subscripts in the figure for brevity. For student $i$ in school $k$, $V_{ik}(z)$ for $z=0,1$ is a function of the student’s fixed pretreatment covariates $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$ and their school membership; $D_{ik}(z)=D_{ik}\left(z,V_{ik}(z)\right)$ for $z=0,1$ is a function of $\mathrm{X}_{ik},\mathrm{U}_{ik},V_{ik}(z)$ and school membership; and $Y_{ik}(z,d)=Y_{ik}\left(z,V_{ik}(z),d\right)$ for $z,d=0,1$ is a function of $\mathrm{X}_{ik},\mathrm{U}_{ik},V_{ik}(\mathrm{z})$ and school membership. Because $D_{ik}(z)$ and $Y_{ik}(z,d)$ are both functions of $V_{ik}(z)$, as shown clearly in Figure 1, $V_{ik}(z)$ is a posttreatment confounder of the causal relationship between $D_{ik}(z)$ and $Y_{ik}(z,d)$ under treatment condition $z$.
@@
Figure 1: Directed Acyclic Graph for Two-Phase Treatment Sequences with Noncompliance.
@@
We specify the simplest form of the theoretical models (i.e., the data generation) for $V_{ik}\left(z\right),D_{ik}\left(z\right)$, and $Y_{ik}\left(z,d\right)$ as the following for $z,d=0,1$. One may expand each model by including site mean-centered $X_{ik}$ as additional predictors. These theoretical models can be extended to the case in which $V_{ik}\left(z\right)$ is multivariate.
$V_{ik}(z)=\alpha_{0ik}+\alpha_{1ik}z;$
$D_{ik}(z)=D_{ik}^{*}(z)+e_{z.ik}^{(D)},\text{ where }D_{ik}^{*}(z)=\beta_{D0ik}+\beta_{D1ik}z+\beta_{Vzk}V_{ik}(z);$
$Y_{ik}(z,d)=\theta_{Y0ik}+\theta_{Vk}V_{ik}(z)+\gamma_{1ik}z+\gamma_{2ik}d+\gamma_{3ik}zd;$ (1)
In the theoretical model for $V_{ik}(z)$, the individual-specific intercept $\alpha_{0ik}$ is equal to $V_{ik}(0)$ explained by $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. The slope $\alpha_{1ik}$ is equal to $V_{ik}(1)-V_{ik}(0)$, which is the individual-specific ITT effect of $Z_{ik}$ on $V_{ik}$ and is also explained by $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$.
Under stochastic noncompliance, let $D_{ik}^{*}(z)=\operatorname{Pr}\left(D_{ik}(z)=1\right)$ represent the student’s probability of attending a small class in Grade 1 after having previously been assigned to kindergarten treatment condition $z$ for $z=0,1$. We assume $0<D_{ik}^{*}(z)<1$ for $z=0,1$. In this model, the intercept $\beta_{D0ik}$ contains the individual-specific element of $D_{ik}^{*}(0)$ that is explained by $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$ but not by $V_{ik}(0)$. The individual-specific ITT effect of $Z_{ik}$ on $D_{ik}^{*}$, which is equal to $D_{ik}^{*}(1)-D_{ik}^{*}(0)$, includes both its direct effect and its indirect effect via $V_{ik}$. The former is captured by $\beta_{D1ik}$ and the latter by $\beta_{V1k}V_{ik}(1)-\beta_{V0k}V_{ik}(0)$, each as a function of $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. In general, the relationship between $V_{ik}(z)$ and $D_{ik}^{*}(z)$ may depend on the Phase I treatment assignment $z$, in which case $\beta_{V1k}\neq\beta_{V0k}$. By definition, $D_{ik}(z)=D_{ik}\left(z,V_{ik}(z)\right)$; and thus $D_{ik}^{*}(z)=D_{ik}^{*}\left(z,V_{ik}(z)\right)$. This model implies three assumptions. (a) Under Phase I treatment assignment $z$ and when $\mathrm{X}_{ik}$, $\mathrm{U}_{ik}$, and $V_{ik}(z)$ are already given, $V_{ik}\left(z^{\prime}\right)$ does not provide additional information about $D_{ik}(z)$ for $z,z^{\prime}=0,1$. This is plausible as a decision about an individual’s Phase II treatment arguably depends on the individual’s observable response to the actual rather than the counterfactual Phase I treatment. (b) The coefficients $\beta_{V1k}$ and $\beta_{V0k}$ are site-specific rather than individual-specific, which we argue is reasonable as individuals within the same site are generally subject to the same decision rules. (c) For simplicity, we specify the theoretical model for $D_{ik}^{*}(z)$ as a linear probability model. The random error term $e_{z.ik}^{(D)}$ exists under stochastic noncompliance. We assume this error term to be independent across individuals under treatment condition $z$ at site $k$, with $E\left\{e_{z.ik}^{(D)}\right\}=0$ for $z=0,1$.
@@
In the theoretical model for $Y_{ik}(z,d)$, the intercept $\theta_{Y0ik}$ captures the individual-specific element of $Y_{ik}(0,0)$ explained by $X_{ik}$ and $U_{ik}$. We use $\gamma_{1ik},\gamma_{2ik}$, and $\gamma_{3ik}$ to denote the components of the cumulative ATE for the individual attributable to the individual’s responses to the Phase I treatment, the Phase II treatment, and their interaction, respectively, when $V_{ik}(z)$ is given. Here $\gamma_{1ik},\gamma_{2ik}$, and $\gamma_{3ik}$ are each a function of $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. The remaining part of the cumulative ATE for the individual is captured by $\theta_{Vk}\left\{V_{ik}(1)-V_{ik}(0)\right\}=\theta_{Vk}\alpha_{1ik}$, which is the indirect effect of $Z_{ik}$ on $Y_{ik}$ transmitted through the individual-specific ITT effect of $Z_{ik}$ on $V_{ik}$ but not through $D_{ik}$. This model implies two assumptions. (a) By definition, $Y_{ik}(z,d)=Y_{ik}\left(z,V_{ik}(z),d\right)$ for $z,d=0,1$, which implies that when $z,d,\mathrm{X}_{ik},\mathrm{U}_{ik}$, and $V_{ik}(z)$ are given, $V_{ik}\left(z^{\prime}\right)$ does not provide additional information about $Y_{ik}(z,d)$. (b) When $z$ and $d$ are given, the association between $V_{ik}$ and $Y_{ik}$ in school $k$, which is represented by $\theta_{Vk}$, is assumed to be linear and not depend on $z$ or $d$.
Although the theoretical models (1) are assumed to be linear and partially additive, analysts may explore alternative model specifications. For instance, to accommodate potential nonlinear relationships between $V$ and $D$ or between $V$ and $Y$, more flexible functional forms can be employed. A continuous $V$ may be discretized into multiple categories with step functions used in the theoretical model for $D$ or $Y$. When $V$ is multivariate, $\theta_{Vk},\beta_{V0k},\beta_{V1k},\alpha_{0k}$, and $\alpha_{1k}$ become vectors.
Aggregating the individual-specific parameters to the site level and using $h$ for numbering, we have that $\alpha_{hk}=E\left(\alpha_{hik}\mid S_{i}=k\right)$ and $\beta_{Dhk}=E\left(\beta_{Dhik}\mid S_{i}=k\right)$ for $h=0,1;\theta_{Y0k}=E\left(\theta_{Y0ik}\mid S_{i}=k\right);$ and $\gamma_{hk}=E\left(\gamma_{hik}\mid S_{i}=k\right)$ for $h=1,2,3$. We will show later that $\beta_{1k}=E\left\{D_{ik}(1)-D_{ik}(0)\mid S_{i}=k\right\}$ represents the school-specific ITT effect of $Z$ on $D$, and that $\beta_{2k}=E\left\{D_{ik}(1)\mid S_{i}=k\right\}$ represents the school-specific compliance rate in Phase II following the experimental condition in Phase I. If school $k$ displays perfect compliance, then we have that $\beta_{1k}=\beta_{2k}=1$. Table 1 provides a glossary of these individual-specific and site-specific parameters. Finally, we define $\alpha_{h}=E\left(\alpha_{hk}\right)$ for $h=0,1;\theta_{V}=E\left(\theta_{Vk}\right);$ and $\gamma_{h}=E\left(\gamma_{hk}\right)$ for $h=1,2,3$.

### v3

@@ 1.1 Relevance
A well-designed intervention often consists of more than one phase. This is because, to prevent fadeouts, the benefit of a treatment in an earlier phase may need to be reinforced through follow-up treatments in the later phases (CunhaHeckman2007). This rationale underpins many sustained treatment strategies in fields such as education, job training, counseling, and medicine for promoting long-term human development and well-being. In evaluating such programs, a research quantity of key scientific interest is the cumulative average treatment effect (ATE) of an intended experimental treatment sequence compared to a control sequence for a target population.
Authorized by the Tennessee state legislature in the 1980s, the Student Teacher Achievement Ratio study (Project STAR) was a state-wide multisite randomized trial designed to evaluate the multi-year cumulative impact of class size reduction on student achievement. Upon kindergarten entry, students within each participating school were randomly assigned to either a small class of 13-17 students (i.e., the experimental group) or a regular class of 22-25 students (i.e., the control group). Students were expected to remain in the class size type initially assigned throughout the study years. At the beginning of each year, teachers within each school were also randomly assigned by the project staff to teach a small class or a regular class (FinnAchilles1990; MostellerLightSachs1995). In this two-year study (kindergarten followed by Grade 1), each year represents a separate phase as students are assessed at the end of a year and are expected to possibly join a different class and meet a different teacher in the coming year. Our application study examines the cumulative ATE of receiving two years of instruction in a small class compared to a regular class in both kindergarten and Grade 1 in student achievement at the end of Grade 1.
@@ 1.2 Methodological Challenges
Noncompliance following the Phase I treatment. Noncompliance often becomes inevitable when an intervention program involves successive phases. In Project STAR, the annual progression in schooling naturally provides an occasion, at the end of each school year, for the school to move some students from one treatment condition to another, even though such decisions might be contrary to the intended research design. Indeed, some students changed their class size types after the kindergarten year (Hanushek1999). Identifying the cumulative ATE is challenging in the presence of noncompliance due to preexisting measured and unmeasured differences between those who complied with the initial treatment assignment and those who did not.
Posttreatment confounding. If an individual does not respond well to a Phase I treatment, it may seem sensible to adjust their treatment in the subsequent phase when possible. In the context of class size interventions, for example, reasons for changing a student’s class size type in Grade 1 are likely tied to the student’s performance in kindergarten. In this case, kindergarten performance is a post-Phase 1 treatment (but pre-Phase II treatment) covariate that may confound the causal relationship between the Grade 1 treatment receipt and the outcome. Importantly, regression-based adjustments for such posttreatment confounding can introduce bias when assessing the effect of the kindergarten (Phase I) treatment on the outcome (Rosenbaum1984).
Non-additive treatment effects. In some cases, continuing the same treatment into a later phase may not only sustain but also amplify the impact of the Phase I treatment. As a result, the benefit of receiving the treatment in two consecutive phases may exceed the sum of the benefits of receiving the treatment in Phase I alone and Phase II alone. In some other cases, a second phase of the same treatment may offer little additional value, such that the combined benefit of receiving two phases of the treatment is less than the sum of the single-phase treatment effects (HongRaudenbush2008). For these reasons, past studies that examined the impact of class size reduction in either kindergarten or Grade 1 in isolation have not addressed the key question about the cumulative ATE of receiving the treatment across both years (Krueger1999).
@@ 1.3 Constraints of Existing Causal Inference Strategies
Inverse-probability-of-treatment weighting (IPTW) is commonly used to address posttreatment confounding when evaluating the cumulative effects of time-varying treatments (HongRaudenbush2008; RobinsHernanBrumback2000) or per-protocol effects in the presence of nonadherence (HernanRobins2017PerProtocol; SmithCoffmanHudgens2021). However, identification using IPTW relies on the sequential ignorability assumption that rarely holds in practice, as noncompliant behavior is often related to unobserved confounders, even after adjusting for observed ones. In prior analyses of the Project STAR data, researchers have employed statistical adjustment for observed pretreatment covariates but failed to account for both posttreatment confounding and unobserved pretreatment confounding (NyeHedgesKonstantopoulos2000).
@@
When applying a standard instrumental variable (IV) strategy to address noncompliance, the analyst may use the initial random treatment assignment as the instrument for the non-random treatment received in a later phase. The IV estimand is defined as the ratio of the ITT effect of the instrument on the outcome to the ITT effect on the treatment received (Bloom1984). This strategy requires the exclusion restriction, which is unlikely to hold in multi-phase interventions. This is because the instrument is expected to have a direct effect on the outcome through actual exposure to the Phase I treatment, regardless of compliance in Phase II. Therefore, the standard IV method is not appropriate for identifying the cumulative ATE in such settings (SwansonLabrecqueHernan2018). Previous studies using Project STAR data have applied the IV method to reduce selection bias associated with noncompliance in year-by-year analyses. However, these approaches do not address the cumulative ATE, as the yearly effects may not be additive (Krueger1999; ShinRaudenbush2011).
@@
Researchers have sought to estimate the average effect of the Phase II treatment on the outcome in the target population by extrapolating from the complier population. This is done by first estimating compliance scores as functions of observed pretreatment covariates that predict both compliant behavior and heterogeneity in treatment effects, and then estimating the average effect of the Phase II treatment within levels of compliance scores (Angrist2010NBER; AronowCarnegie2013; EsterlingNebloLazer2011; WangRobinsRichardson2017). These strategies nonetheless rely on a form of the strong ignorability assumption that is often unrealistic. Specifically, unobserved pretreatment confounders as well as observed and unobserved time-varying confounders may not be captured by the estimated compliance scores, which undermines the validity of the identification.
@@
The multiple-site multiple-mediator IV (MSMM-IV) strategy (BloomUntermanZhuReardon2020; DuncanMorrisRodrigues2011; KlingLiebmanKatz2007; NomiRaudenbushSmith2021; RaudenbushReardonNomi2012; ReardonRaudenbush2013), developed to identify the causal effects of multiple parallel mediators in a multisite trial, similarly requires the exclusion restriction and overlooks non-additive cumulative treatment effects. This strategy generally does not allow for time-varying confounders of the mediator-outcome relationships that may affect the outcome both directly and indirectly (see ReardonRaudenbush2013 for a discussion of a possible relaxation of this constraint).
@@ 1.4 Proposed Multisite Two-Phase Treatment IV Strategy
When using the initial random treatment assignment as an instrument to address complications arising from noncompliance, the well-known principal stratification framework (FrangakisRubin2002) assumes deterministic noncompliance. That is, each individual is classified as either a complier or a noncomplier, based on the notion that compliance is pre-determined and fixed. Under this framework, the IV estimand identifies the Complier Average Treatment Effect (CATE), also known as the Local Average Treatment Effect (LATE) (AngristImbensRubin1996).
However, as SmallTanEtAl2017 argue, this deterministic view may not align with reality, especially when compliant behavior is influenced by stochastic events. In such cases, the IV estimand instead identifies a Weighted Average Treatment Effect (WATE), a quantity that often does not directly correspond to the target population of the study. In our application, we similarly conceptualize noncompliance as stochastic. For the primary target population in Project STAR, due to unforeseen opportunities and constraints, whether a student would change class type in Grade 1 is probable rather than pre-determined. Therefore, we argue that the concepts of CATE or LATE may not be applicable while the concept of cumulative ATE of attending a small versus regular class over consecutive years is relevant.
Our proposed approach is specific to multisite randomized trials. But it may also be applicable to settings analogous to two-phase treatment designs. For example, HeckmanSmithTaber1998 demonstrated that intervention programs may affect not only experimental participants who receive the full treatment but also those who receive a partial “dose” before dropping out. When the effect of partial treatment is nonzero, the IV estimator fails to identify the average treatment effect for the fully treated due to the violation of the exclusion restriction. In such a case, Heckman and colleagues argued that an economically relevant causal parameter is the average treatment effect that would result if the entire experimental group would receive the full treatment. This parameter aligns with our concept of cumulative ATE, with individuals receiving partial treatment prior to dropout analogous to those who switch from the experimental condition to the control condition after receiving a Phase-I treatment.
@@ 2.1 Notation and the Causal Estimand
Let $Z_{ik}$ be a binary indicator for the Phase I treatment that takes value $z=1$ if individual $i$ at site $k$ is assigned to the experimental condition and 0 if assigned to the control condition at the baseline. In Project STAR, due to full compliance in kindergarten, the treatment assigned is equivalent to the treatment received for each student in Phase I. Let $D_{ik}$ be a binary indicator for the Phase II treatment receipt that takes value $d=1$ if the individual receives the experimental treatment and 0 if receiving the control treatment. The individual’s potential outcome by the end of Phase II can be represented as a function of the initial randomized treatment assignment $z$ or a function of the two-phase treatment sequence $z$ and $d$. It is typically denoted as $Y_{ik}(z)$ in the former case and $Y_{ik}(z,d)$ in the latter case for $z=0,1$ and $d=0,1$ under the stable unit treatment value assumption (SUTVA). In other words, we assume that there is a single version of each treatment and that there is no interference among individuals between or within sites. Similarly, we may use $V_{ik}(z)$ to denote the individual’s intermediate potential outcome by the end of Phase I and use $D_{ik}(z)$ to denote the individual’s Phase-II treatment receipt each as a function of the individual’s Phase-I treatment assignment. We use $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$ to denote the individual’s observed and unobserved vectors of baseline covariates, respectively. We use $S_{i}=k$ to indicate that individual $i$ is a member of site $k$.
Corresponding to a research question that initially motivated Project STAR, the causal effect of attending a small class versus a regular class in both kindergarten and Grade 1 for student $i$ in school $k$ is $Y_{ik}(1,1)-Y_{ik}(0,0)$. The cumulative ATE of the two-phase treatment, as defined below, represents the average of site-specific treatment effects for a population of sites where the inner expectation is taken over the population of individuals within site $k$ and the outer expectation is taken over all the sites:
@@ 2.2 Theoretical Model
Figure 1 illustrates the conceptual relationships between $X,U,Z,V,D$, and $Y$. We have omitted the subscripts in the figure for brevity. For individual $i$ in site $k$, $V_{ik}(z)$ for $z=0,1$ is a function of the individual’s fixed pretreatment covariates $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$ and site membership; and $Y_{ik}(z,d)=Y_{ik}\left(z,V_{ik}(z),d\right)$ for $z,d=0,1$ is a function of $\mathrm{X}_{ik},\mathrm{U}_{ik},V_{ik}(\mathrm{z})$ and site membership. Importantly, $D_{ik}(z)$ for $z=0,1$ is subject to the selection associated with $\mathrm{X}_{ik},\mathrm{U}_{ik}$, and $V_{ik}(z)$ within a site. Because $D_{ik}(z)$ and $Y_{ik}(z,d)$ are both functions of $V_{ik}(z)$, as shown clearly in Figure 1, $V_{ik}(z)$ is a posttreatment confounder of the causal relationship between $D_{ik}(z)$ and $Y_{ik}(z,d)$ under treatment condition $z$.
@@
Figure 1: Directed Acyclic Graph for Two-Phase Treatment Sequences with Noncompliance. $Z$ and $D$ respectively denote the Phase I treatment and Phase II treatment receipt. $V$ and $Y$ respectively stand for the outcome by the end of Phase I and that by the end of Phase II. $X$ and $U$ denote the observed and unobserved vectors of baseline covariates, respectively.
Let $\beta_{0ik}=D_{ik}(0)$ denote an individual’s Phase-II treatment after being assigned to the control condition in Phase I; let $\beta_{1ik}=D_{ik}(1)-D_{ik}(0)$ denote the individual-specific ITT effect of $Z_{ik}$ on $D_{ik}$; and let $\beta_{2ik}=D_{ik}(1)$ denote the individual’s Phase-II treatment after being assigned to the experimental condition in Phase I. Here $\beta_{0ik}$, $\beta_{1ik}$, and $\beta_{2ik}$ are each a function of $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. We then define $\beta_{0k}=E\left\{\beta_{0ik}\mid S_{i}=k\right\}$ as the site-specific noncompliance rate in Phase II following the assignment to the control condition in Phase I, $\beta_{1k}=E\left\{\beta_{1ik}\mid S_{i}=k\right\}$ as the site-specific ITT effect of $Z$ on $D$, and $\beta_{2k}=E\left\{\beta_{2ik}\mid S_{i}=k\right\}$ as the site-specific compliance rate in Phase II following the assignment to the experimental condition in Phase I. If site $k$ displays perfect compliance, then we have that $\beta_{0k}=0$ and $\beta_{1k}=\beta_{2k}=1$.
In addition, let $\alpha_{0ik}=V_{ik}(0)$ denote an individual’s Phase-I outcome after being assigned to the control condition in Phase I, which is explained by $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. Let $\alpha_{1ik}=V_{ik}(1)-V_{ik}(0)$, which is the individual-specific ITT effect of $Z_{ik}$ on $V_{ik}$ and is also explained by $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. $V$ can be either continuous or discrete. When $V$ is multi-categorical or multivariate, $\alpha_{0ik}$ and $\alpha_{1ik}$ become vectors.
@@
We now specify the simplest form of the theoretical model (i.e., the data generation) for $Y_{ik}\left(z,d\right)$ as the following for $z,d=0,1$. One may expand this model by including site mean-centered $X_{ik}$ as additional predictors. The theoretical model can be extended to the case in which $V_{ik}\left(z\right)$ is multivariate.
$Y_{ik}(z,d)=\theta_{Y0ik}+\theta_{Vik}V_{ik}(z)+\gamma_{1ik}z+\gamma_{2ik}d+\gamma_{3ik}zd;$ (1)
@@
In the theoretical model for $Y_{ik}(z,d)$, the intercept $\theta_{Y0ik}$ captures the individual-specific element of $Y_{ik}(0,0)$ explained by $X_{ik}$ and $U_{ik}$. We use $\gamma_{1ik},\gamma_{2ik}$, and $\gamma_{3ik}$ to denote the components of the cumulative ATE for the individual attributable to the individual’s responses to the Phase I treatment, the Phase II treatment, and their interaction, respectively, when $V_{ik}(z)$ is given. Here $\gamma_{1ik},\gamma_{2ik}$, and $\gamma_{3ik}$ are each a function of $\mathrm{X}_{ik}$ and $\mathrm{U}_{ik}$. The remaining part of the cumulative ATE for the individual is captured by $\theta_{Vik}\left\{V_{ik}(1)-V_{ik}(0)\right\}=\theta_{Vik}\alpha_{1ik}$, which is the indirect effect of $Z_{ik}$ on $Y_{ik}$ transmitted through the individual-specific ITT effect of $Z_{ik}$ on $V_{ik}$ but not through $D_{ik}$. This model implies two assumptions. (a) By definition, $Y_{ik}(z,d)=Y_{ik}\left(z,V_{ik}(z),d\right)$ for $z,d=0,1$, which implies that when $z,d,\mathrm{X}_{ik},\mathrm{U}_{ik}$, and $V_{ik}(z)$ are given, $V_{ik}\left(z^{\prime}\right)$ does not provide additional information about $Y_{ik}(z,d)$. (b) When $z$ and $d$ are given, the association between $V_{ik}$ and $Y_{ik}$, which is represented by $\theta_{Vik}$, is assumed to be linear and not depend on $z$ or $d$. Although this theoretical model (1) is assumed to be linear and partially additive, analysts may explore alternative model specifications. For instance, to accommodate potential nonlinear relationships between $V$ and $Y$, more flexible functional forms can be employed.
Aggregating the individual-specific parameters to the site level and using $h$ for numbering, we have that $\alpha_{hk}=E\left(\alpha_{hik}\mid S_{i}=k\right)$ for $h=0,1;\theta_{Y0k}=E\left(\theta_{Y0ik}\mid S_{i}=k\right);$ and $\gamma_{hk}=E\left(\gamma_{hik}\mid S_{i}=k\right)$ for $h=1,2,3$.
Table 1 provides a glossary of these individual-specific and site-specific parameters.
Finally, we define $\theta_{V}=E\left(\theta_{Vk}\right)$ and $\gamma_{h}=E\left(\gamma_{hk}\right)$ for $h=1,2,3$.